OpenStudy (anonymous):

\[Evaluate \int\limits_{1}^{2} \frac{2^{x}-3^{x}}{3^{^{x}}}dx\]

OpenStudy (experimentx):

\[ \frac{2^x}{3^x} - 1 = \left ( e^{\ln{2/3}x}\right ) - 1\]